# Asymmetric Errors

### v1

@@ 2.3 Combination of pdf results
As discussed in Section 1.1, the use of likelihoods to combine results is a very standard procedure and their use to produce a combined error is less mainstream, whereas for pdfs the case is reversed. In fact, for pdfs the combination of results can be considered as an instance of the combination of errors. The measurements are all of the same quantity and combined by taking an average, rather than distinct quantities combined with some general function.
If several estimates $\hat{r}_{i}$ with pdfs of variances $V_{i}$ are obtained by different experiments for the same result $r$, then an obvious way to form a combined result is to take a weighted sum:
@@
Input Combined result using model
$r_{1}$ $r_{2}$ Dimidiated Gaussian Distorted Gaussian Double Cubic Gaussian Edgeworth Expansion Fechner Distribution
${32.571}^{+7.571}_{-6.571}$ ${18.429}^{+7.571}_{-6.571}$ ${25.700}^{+5.252}_{-4.752}$ $4.763{25.750}^{++}_{-5.262}$ 4.763 ${25.699}^{+5.254}_{-4.754}$ ${25.749}^{+5.251}_{-4.749}$ ${25.712}^{+5.252}_{-4.7492}$
Johnson System Log Normal QVW Gaussian Railway Gaussian Skew Normal
${25.742}^{+5.333}_{-4.809}$ ${25.748}^{+5.282}_{-4.776}$ ${25.707}^{+5.255}_{-4.755}$ ${25.749}^{+5.261}_{-4.765}$ ${25.741}^{+5.255}_{-4.777}$
Symmetric Beta Gaussian(1,10) Symmetric Beta Gaussian(1,15) Symmetric Beta Gaussian(1,20) Symmetric Beta Gaussian(1,25) Symmetric Beta Gaussian(1,30)
${25.700}^{+5.254}_{-4.755}$ ${25.717}^{+5.256}_{-4.757}$ ${25.730}^{+5.258}_{-4.759}$ ${25.737}^{+5.259}_{-4.760}$ ${25.741}^{+5.260}_{-4.761}$
Symmetric Beta Gaussian(2,10) Symmetric Beta Gaussian(2,15) Symmetric Beta Gaussian(2,20) Symmetric Beta Gaussian(2,25) Symmetric Beta Gaussian(2,30)
${25.697}^{+5.254}_{-4.754}$ ${25.708}^{+5.255}_{-4.756}$ ${25.720}^{+5.256}_{-4.757}$ ${25.728}^{+5.258}_{-4.759}$ ${25.734}^{+5.259}_{-4.759}$
Symmetric Beta Gaussian(3,10) Symmetric Beta Gaussian(3,15) Symmetric Beta Gaussian(3,20) Symmetric Beta Gaussian(3,25) Symmetric Beta Gaussian(3,30)
${25.696}^{+5.253}_{-4.754}$ ${25.704}^{+5.254}_{-4.755}$ ${25.714}^{+5.256}_{-4.756}$ ${25.722}^{+5.257}_{-4.758}$ ${25.729}^{+5.258}_{-4.759}$
Symmetric Beta Gaussian(4,10) Symmetric Beta Gaussian(4,15) Symmetric Beta Gaussian(4,20) Symmetric Beta Gaussian(4,25) Symmetric Beta Gaussian(4,30)
${25.695}^{+5.253}_{-4.754}$ ${25.701}^{+5.254}_{-4.755}$ ${25.710}^{+5.255}_{-4.756}$ ${25.718}^{+5.256}_{-4.757}$ ${25.725}^{+5.257}_{-4.758}$
$r_{1}$ $r_{2}$ Dimidiated Gaussian Distorted Gaussian Double Cubic Gaussian Edgeworth Expansion Fechner Distribution
${41.142}^{+7.571}_{-6.571}$ ${12.878}^{+7.571}_{-6.571}$ ${27.200}^{+5.252}_{-4.752}$ ${27.250}^{+5.262}_{-4.763}$ ${27.199}^{+5.254}_{-4.754}$ ${27.249}^{+5.251}_{-4.74}$ ${27.212}^{+5.252}_{-4.749}$
Johnson System Log Normal QVW Gaussian Railway Gaussian Skew Normal
${27.242}^{+5.333}_{-4.809}$ ${27.248}^{+5.282}_{-4.776}$ ${27.207}^{+5.255}_{-4.755}$ ${27.249}^{+5.261}_{-4.765}$ ${27.241}^{+5.255}_{-4.777}$
Symmetric Beta Gaussian(1,10) Symmetric Beta Gaussian(1,15) Symmetric Beta Gaussian(1,20) Symmetric Beta Gaussian(1,25) Symmetric Beta Gaussian(1,30)
${27.200}^{+5.254}_{-4.755}$ ${27.217}^{+5.256}_{-4.757}$ ${27.230}^{+5.258}_{-4.759}$ ${27.237}^{+5.259}_{-4.760}$ ${27.241}^{+5.260}_{-4.761}$
Symmetric Beta Gaussian(2,10) Symmetric Beta Gaussian(2,15) Symmetric Beta Gaussian(2,20) Symmetric Beta Gaussian(2,25) Symmetric Beta Gaussian(2,30)
${27.197}^{+5.254}_{-4.754}$ ${27.208}^{+5.255}_{-4.756}$ ${27.220}^{+5.256}_{-4.757}$ ${27.228}^{+5.258}_{-4.759}$ ${27.234}^{+5.259}_{-4.759}$
Symmetric Beta Gaussian(3,10) Symmetric Beta Gaussian(3,15) Symmetric Beta Gaussian(3,20) Symmetric Beta Gaussian(3,25) Symmetric Beta Gaussian(3,30)
${27.196}^{+5.253}_{-4.754}$ ${27.204}^{+5.254}_{-4.755}$ ${27.214}^{+5.256}_{-4.756}$ ${27.222}^{+5.257}_{-4.758}$ ${27.229}^{+5.258}_{-4.759}$
Symmetric Beta Gaussian(4,10) Symmetric Beta Gaussian(4,15) Symmetric Beta Gaussian(4,20) Symmetric Beta Gaussian(4,25) Symmetric Beta Gaussian(4,30)
${27.195}^{+5.253}_{-4.754}$ ${27.201}^{+5.254}_{-4.755}$ ${27.210}^{+5.255}_{-4.756}$ ${27.218}^{+5.256}_{-4.757}$ ${27.225}^{+5.257}_{-4.758}$
Table 3: Examples of combination of errors with pdfs using different models - ARB
@@
Model $V_{pred}$ $\frac{V_{pred}-V_{MC}}{V_{MC}}$ $\gamma_{pred}$
Dimidiated Gaussian 25.045 -0.0096 14.967
Distorted Gaussian 25.250 -0.0015 37.750
Double Cubic Gaussian 25.083 -0.0081 20.765
Edgeworth Expansion 25.000 -0.0114 37.751
Fechner Distribution 25.016 -0.0107 25.157
Johnson System 26.555 0.0501 44.119
Log Normal 25.593 0.0121 39.367
QVW Gaussian 25.103 -0.0073 23.277
Railway Gaussian 25.249 -0.0015 36.867
Skew Normal 25.583 0.0117 33.617
Symmetric Beta Gaussian(1,10) 25.090 -0.0078 21.804
Symmetric Beta Gaussian(1,30) 25.206 -0.0032 34.141
Symmetric Beta Gaussian(4,10) 25.071 -0.0086 19.007
Symmetric Beta Gaussian(4,30) 25.150 -0.0055 28.741
Table 4: Comparison of quoted quantities and toy Monte Carlo results (based upon 1,000,000 toys). ARB: Need to think about this table…$\mu_{pred}$ is based on a single measurement (random number), while $\mu_{MC}$ is the MC mean of the distribution of the combined median. Check whether $V_{pred}$ and $\gamma_{pred}$ only depend on $\sigma_{+}$ and $\sigma_{-}$ for all models; in case, simplify the table.
This framework may be unrealistic, in that it assumes that the (asymmetric) errors are independent of the measurement. The quoted errors on the ‘input’ column in Table 3 are all the same, because the true mean of $x$ is known. This will be complicated by additional information from the experiments which does not appear in the quoted values and errors. This is considered in detail by Schmelling[Schmelling] using a second order approximation (the “distorted Gaussian” in this work). Rather than the ideal case where all true quantities are known, he considers the more realistic case where the errors are taken from measurements (as happens when $\sqrt{N}_{observed}$ errors are used for Poisson statistics), and shows that as well as the bias from the difference between median and mean there is a bias from the use of an estimator $\hat{V}$ for the variance, and a further significant bias from the correlation between $V$ and $\hat{r}$. If the parameter of interest is, for example, the square of a Gaussian quantity, then upward fluctuations will be ascribed larger errors, and thus smaller weights.
@@
For a given measurement ${x_{i}}^{+{\sigma^{+}_{i}}}_{-{\sigma^{-}_{i}}}$ we can express its compatibility with some quoted $x_{c}$ using the p-value or its equivalent expression in terms of sigmas. With the familiar Gaussian, we would say that a measurement of $12.7\pm 0.1$ was 5 sigma away from some nominal value of $12.2$, and accordingly treat it with strong reservation. A similar approach for asymmetric errors will use ${\sigma^{+}_{i}}$ and ${\sigma^{-}_{i}}$, but care must be taken over the direction of the deviation. Using the simple dimidiated model, a value of ${12.7}^{+0.1}_{-0.2}$ is 2.5 sigma away from a proposed true value of $12.2$, not 5 sigma, as if a true 12.2 gives a measured 12.7 this is an upward fluctuation. For other models the arithmetic is not so simple, but the p-value can readily be determined from the pdf.
@@ 3.1 Modelling non-parabolic likelihoods
For the Gaussian the log likelihood function $\ln L$ is a parabola with two parameters, the location and scale. (The parameter giving the value at the peak is irrelevant.) For an asymmetric form a reasonable guess at the full likelihood function will be some curve with three parameters, approximately parabolic, for which the maximum likelihood peak occurs at some $a=\hat{a}$ and $\Delta\ln L=-{1\over 2}$ at $a=\hat{a}+{\sigma^{+}}$ and $a=\hat{a}-{\sigma^{-}}$. Thus the curve must go through 3 points, having a maximum at the middle one. The 3 parameters will be obtainable from the values of $\hat{a},{\sigma^{+}}$ and ${\sigma^{-}}$, plus an irrelevant term giving the actual value at the peak. It should also have a ‘reasonable’ behaviour elsewhere; in particular it should go to $\infty$- ∞ at large positive and negative $a$. Provision of such models is somewhat easier for log likelihoods than for pdfs, as there are only two sets of parameters deployed: these $\hat{a},{\sigma^{+}},{\sigma^{-}}$ and the 3 parameters of the specific model. The software must convert between the two, but there is no third set corresponding to the moments of the pdfs.
Two models which are simple to use and turn out to be generally applicable arise from the suggestion by Bartlett that a likelihood function is described by a Gaussian whose width varies as a function of the argument $a$ [bartlett1, bartlett2, deltalnL]. With such an assumption one might suppose that the variation is linear, either for the standard deviation (‘linear sigma model’) or the variance (‘linear variance model’).
@@
Figure 8:
Using the linear sigma and linear variance models of the log likelihood arising from ${5.0}^{+1.1}_{-0.9}$ (left) and ${5.0}^{+1.5}_{-0.5}$ (right)
@@ 3.1.1 Comparison of models and recommendations
Linear variance $V(a)=V+aV^{\prime}$ Good general-purpose use
Finite domain
Cubic $-{1\over 2}(\alpha a^{2}+\beta a^{3})$ Badly behaved, $\ln L$ turns over
Broken parabola $-{1\over 2}{a^{2}\over\sigma_{\pm}^{2}}$ Simple but $L^{\prime\prime}$ discontinuity at peak
Constrained quartic $-{1\over 2}(\alpha+\beta a)^{2}$ quartic constrained a single peak
Molded quartic $-{1\over 2}(\alpha a^{4}+\beta a^{3}+\gamma a^{2})$ quartic is best match to broken
parabola in central region
Matched quintic Quintic centrally, $\sigma_{\pm}$ outside Quintic coeffts from matching
$L$ and $L^{\prime\prime}$ at $\sigma_{\pm}$
Interpolated seventh 7th order centrally, $\sigma_{\pm}$ outside Coeffts from matching $L,L^{\prime},L^{\prime\prime}$
Simple double Central quartics, quadratics outside Coeffts from matching
quartic $\sigma_{0}=\sqrt{{\sigma^{-}}{\sigma^{+}}}$
Molded double Central quartics, quadratics outside Coeffts from matching
quartic $\sigma_{0}=\sqrt{({\sigma^{-}}^{4}+{\sigma^{+}}^{4})/({\sigma^{-}}^{2}+{\sigma^{+}}^{2}}$
Double quintic Central quintics, quadratics outside
Conservative spline Cubic between spline points, Spline points need to be chosen
quadratics outside in central region
Logarithmic $-{1\over 2}({\ln(1+\gamma a)\over\ln\beta})^{2}$ Parabola subject to scaled
expansion. Limited domain.
Generalised Poisson $-\alpha a+{\cal N}(1+{\alpha a\over{\cal N}})$ Poisson, but ${\cal N}$ need not be integer
Linear sigma log $\sigma=e^{\alpha Q(a)+\beta}$ $Q(a)$ is the cdf of the equivalent
Fechner distribution
Double cubic log $\sigma=e^{cubics}$ centrally, $\sigma_{\pm}$ outside Different cubics for+ve and -ve $a$.
Quintic sigma log $\sigma=e^{quintic}$ centrally, $\sigma_{\sigma^{+}}$ outside Quintic by matching $L,L^{\prime},L^{\prime\prime}$
PDG $\sigma(a)=\sigma+a\sigma^{\prime}$ or $\sigma_{\pm}$ Linear $\sigma$ centrally,
${\sigma^{+}}$ and ${\sigma^{-}}$ above & below
Edgeworth Log of Eq. (A.8) Numerically messy
Skew normal Log of Eq (44) Numerically messy
@@
As an illustration we take the result $a={5.0000}^{+2.5811}_{-1.9159}$, which would be reported from a Poisson measurement of 5 events. This is shown in Figure 9. The true likelihood is $L(a)=e^{-a}{a^{5}\over 5!}$, and we compare this with the results of the likelihoods given by the models. All curves, by construction, peak at $a=5$ and go through the $\Delta\ln L=\pm{1\over 2}$ points, and their interpolations in the central region are very similar. At larger values their behaviours are very different. The cubic turns over, which is unacceptable. The quartic provides a better match to the original Poisson likelihood, especially on the negative side. The logarithmic form does fairly well, the Poisson does perfectly, as it should in this particular example. The two Bartlett forms both do well, the linear variance is somewhat better than the linear sigma. The PDG form does badly, as it assumes a Gaussian behaviour outside the central interval. The Edgeworth form does badly at both small and large values, and is troubled by logarithms of negative numbers at large values and/or asymmetries. The skew normal is not as bad as Edgeworth’s approximation, but it does not do very well. (These comparisons assume the fundamental $a={5.0000}^{+2.5811}_{-1.9159}$ measurement comes from a Poisson measurement: if it does not, then conclusions might be different.)
@@
As another example, this time with a negative skewness, we take the logarithm of a parameter with a likelihood described by a Gaussian with $\mu=8.0$ and $\sigma=3.0$ This is shown in Figure 10. The reported result is $a={2.08}^{+0.32}_{-0.47}$. Again the cubic does predictably badly, and the quartic surprisingly well. The logarithmic does quite well, and the Poisson is not good. The PDG form is again bad, but the Bartlett forms are good: this time the linear sigma does a bit better than the linear variance. The Edgeworth form does badly and the skew normal form does not do particularly well.
Figure 11: The central regions for Figures 9 and 10.

### v2

@@ 2.4 Combination of pdf results
As discussed in Section 1.1, the use of likelihoods to combine results is a very standard procedure and their use to produce a combined error is less mainstream, whereas for pdfs the case is reversed. In fact, for pdfs the combination of results can be considered as an instance of the combination of errors. The measurements are all of the same quantity and combined by taking a weighted meanrather than distinct quantities combined with some general function.
If several independentestimates $\hat{r}_{i}$ with pdfs of variances $V_{i}$ are obtained by different experiments for the same result $r$, then an obvious way to form a combined result is to take a weighted sum:
@@
Input Combined result using model
$r_{1}$ $r_{2}$ Dimidiated Gaussian Distorted Gaussian
${32.571}^{+7.571}_{-6.571}$ ${18.429}^{+7.571}_{-6.571}$ ${25.700}^{+5.252}_{-4.752}$ ${25.750}^{+5.262}_{-4.763}$
Railway Gaussian Double Cubic Gaussian
${25.749}^{+5.261}_{-4.765}$ ${25.699}^{+5.254}_{-4.754}$
QVW Gaussian Fechner Distribution
${25.707}^{+5.255}_{-4.755}$ ${25.712}^{+5.252}_{-4.749}$
Edgeworth expansion Skew normal
${25.749}^{+5.251}_{-4.749}$ ${25.741}^{+5.255}_{-4.777}$
Johnson System Log normal
${25.742}^{+5.333}_{-4.809}$ ${25.748}^{+5.282}_{-4.776}$
Sym. beta Gaussian(1,10) Sym. beta Gaussian(1,30)
${25.700}^{+5.254}_{-4.755}$ ${25.741}^{+5.260}_{-4.761}$
Sym. beta Gaussian(4,10) Sym. beta Gaussian(4,30)
${25.695}^{+5.253}_{-4.754}$ ${25.725}^{+5.257}_{-4.758}$
$r_{1}$ $r_{2}$ Dimidiated Gaussian Distorted Gaussian
${41.142}^{+7.571}_{-6.571}$ ${12.878}^{+7.571}_{-6.571}$ ${27.200}^{+5.252}_{-4.752}$ ${27.250}^{+5.262}_{-4.763}$
Railway Gaussian Double Cubic Gaussian
${27.249}^{+5.261}_{-4.765}$ ${27.199}^{+5.254}_{-4.754}$
QVW Gaussian Fechner Distribution
${27.207}^{+5.255}_{-4.755}$ ${27.212}^{+5.252}_{-4.749}$
Edgeworth expansion Skew normal
${27.249}^{+5.251}_{-4.74}$ ${27.241}^{+5.255}_{-4.777}$
Johnson System Log normal
${27.242}^{+5.333}_{-4.809}$ ${27.248}^{+5.282}_{-4.776}$
Sym. beta Gaussian(1,10) Sym. beta Gaussian(1,30)
${27.200}^{+5.254}_{-4.755}$ ${27.241}^{+5.260}_{-4.761}$
Sym. beta Gaussian(4,10) Sym. beta Gaussian(4,30)
${27.195}^{+5.253}_{-4.754}$ ${27.225}^{+5.257}_{-4.758}$
Table 4: Examples of combination of errors with pdfs using different models.
@@
Model $V_{pred}$ $\frac{V_{pred}-V_{MC}}{V_{MC}}$ $\gamma_{pred}$ $\frac{\gamma_{pred}-\gamma_{MC}}{\gamma_{MC}}$
Dimidiated Gaussian 25.045 -0.0096 14.967 -0.610
Distorted Gaussian 25.250 -0.0015 37.750 -0.015
Railway Gaussian 25.249 -0.0015 36.867 -0.038
Double cubic Gaussian 25.083 -0.0081 20.765 -0.458
Symmetric beta Gaussian(1,10) 25.090 -0.0078 21.804 -0.431
Symmetric beta Gaussian(1,30) 25.206 -0.0032 34.141 -0.109
Symmetric beta Gaussian(4,10) 25.071 -0.0086 19.007 -0.504
Symmetric beta Gaussian(4,30) 25.150 -0.0055 28.741 -0.250
QVW Gaussian 25.103 -0.0073 23.277 -0.393
Fechner distribution 25.016 -0.0107 25.157 -0.344
Edgeworth expansion 25.000 -0.0114 37.751 -0.015
Skew normal 25.583 0.0117 33.617 -0.123
Johnson system 26.555 0.0501 44.119 0.151
Log normal 25.593 0.0121 39.367 0.027
Table 5: Comparison of quoted quantities and toy Monte Carlo results based upon $10^{6}$ toys. The toy results are not identical for all models but agree to five significant digits, $V_{MC}=25.288$ and $\gamma_{MC}=38.388$.
This framework may be unrealistic, in that it assumes that the (asymmetric) errors are independent of the measurement. The quoted errors on the ‘input’ column in Table 4 are all the same, because the true mean of $x$ is known. This will be complicated by additional information from the experiments which does not appear in the quoted values and errors. This is considered in detail by Schmelling [6] using a second order approximation (the “distorted Gaussian” in this work). Rather than the ideal case where all true quantities are known, he considers the more realistic case where the errors are taken from measurements (as happens when $\sqrt{N}_{observed}$ errors are used for Poisson statistics), and shows that as well as the bias from the difference between median and mean there is a bias from the use of an estimator $\hat{V}$ for the variance, and a further significant bias from the correlation between $V$ and $\hat{r}$. If the parameter of interest is, for example, the square of a Gaussian quantity, then upward fluctuations will be ascribed larger errors, and thus smaller weights.
@@
For a given measurement ${x_{i}}^{+{\sigma^{+}_{i}}}_{-{\sigma^{-}_{i}}}$ we can express its compatibility with some quoted $x_{c}$ using the p-value or its equivalent expression in terms of sigmas. With the familiar Gaussian, we would say that a measurement of $12.7\pm 0.1$ was 5 sigma away from some nominal value of $12.2$, and accordingly treat it with strong reservation. A similar approach for asymmetric errors will use ${\sigma^{+}_{i}}$ and ${\sigma^{-}_{i}}$, but care must be taken over the direction of the deviation. Suppose the proposed true value is $12.2$ and the measurement is ${12.7}^{+0.1}_{-0.2}$. Using the dimidiated model, this deviation is a 5 sigma, not 2.5 sigma, as the quoted measurement implies that a process which may go up a little or down a lot has given a result of 12.7. For other models the arithmetic is not so simple, but the p-value can readily be determined from the pdf.
@@ 3.1 Modelling non-parabolic likelihoods
For the Gaussian the log likelihood function $\ln L$ is a parabola with two parameters, the location and scale. (The parameter giving the value at the peak is irrelevant.) For an asymmetric form a reasonable guess at the full likelihood function will be some curve with three parameters, approximately parabolic, for which the maximum likelihood peak occurs at some $a=\hat{a}$ and $\Delta\ln L=-{1\over 2}$ at $a=\hat{a}+{\sigma^{+}}$ and $a=\hat{a}-{\sigma^{-}}$. Thus the curve must go through 3 points, having a maximum at the middle one. The 3 parameters will be obtainable from the values of $\hat{a},{\sigma^{+}}$ and ${\sigma^{-}}$. It should also have a ‘reasonable’ behaviour elsewhere; in particular it should go to $\infty$- ∞ at large positive and negative $a$. Provision of such models is somewhat easier for log likelihoods than for pdfs, as there are only two sets of parameters deployed: these $\hat{a},{\sigma^{+}},{\sigma^{-}}$ and the 3 parameters of the specific model. The software must convert between the two, but there is no third set corresponding to the moments of the pdfs.
Two models which are simple to use and turn out to be generally applicable arise from the suggestion by Bartlett that a likelihood function is described by a Gaussian whose width varies as a function of the argument $a$ [16, 17, 13]. With such an assumption one might suppose that the variation is linear, either for the standard deviation (‘linear sigma model’) or the variance (‘linear variance model’).
@@
Figure 9: Using the linear sigma and linear variance models of the log likelihood arising from ${5.0}^{+1.1}_{-0.9}$ (left) and ${5.0}^{+0.5}_{-1.5}$(right).
@@
If any information is provided, or can be deduced, about the form of the likelihood then that should be used to guide the choice of model – there is no general ‘best model’, but some models do better for particular classes of result. By the same token, if a result with an asymmetric error is quoted any further available information should be provided. If all the information available to a meta-analyst is the 3 numbers of the reported measurement, then the wide range of available models can potentially give a wide range of outcomes.
@@ 3.1.1 Comparison of models and recommendations
Cubic $-{1\over 2}(\alpha a^{2}+\beta a^{3})$ Badly behaved, $\ln L$ turns over.
Broken parabola $-{1\over 2}{a^{2}\over\sigma_{\pm}^{2}}$ Simple, but $L^{\prime\prime}$ discontinuity at peak
Constrained quartic $-{1\over 2}(\alpha+\beta a)^{2}$ Quartic constrained at single peak
Molded quartic $-{1\over 2}(\alpha a^{4}+\beta a^{3}+\gamma a^{2})$ Quartic is best match to broken parabola in central region.
Matched quintic Quintic centrally, $\sigma_{\pm}$ outside Quintic coeffts from matching $L$ and $L^{\prime\prime}$ at $\sigma_{\pm}$
Interpolated 7th degree 7th order centrally, $\sigma_{\pm}$ outside Coeffts from matching $L,L^{\prime},L^{\prime\prime}$
Simple double quartic Central quartics, quadratics outside $\sigma_{0}=\sqrt{{\sigma^{-}}{\sigma^{+}}}$ Coeffts from matching
Molded double quartic Central quartics, quadratics outside $\sigma_{0}=\sqrt{({\sigma^{-}}^{4}+{\sigma^{+}}^{4})/({\sigma^{-}}^{2}+{\sigma^{+}}^{2}}$ Coeffts from matching
Simple double quintic Central quintics, quadratics outside
Conservative spline Cubic between spline points, quadratics outside Spline points need to be chosen in central region.
Logarithmic $-{1\over 2}({\ln(1+\gamma a)\over\ln\beta})^{2}$ Parabola subject to scaled expansion. Limited domain
Generalised Poisson $-\alpha a+{\cal N}(1+{\alpha a\over{\cal N}})$ Poisson, but ${\cal N}$ need not be integer.
@@
Linear variance $V(a)=V+aV^{\prime}$ Good general-purpose use. Finite domain
Linear log sigma $\sigma=e^{\alpha Q(a)+\beta}$ $Q(a)$ is the cdf of the equivalent Fechner distribution.
Double cubic log sigma $\sigma=e^{cubics}$ centrally, $\sigma_{\pm}$ outside Different cubics for $a\leq 0$ and $a\geq 0$
Quintic log sigma $\sigma=e^{quintic}$ centrally, $\sigma_{\sigma^{+}}$ outside Quintic by matching $L,L^{\prime},L^{\prime\prime}$
PDG method $\sigma(a)=\sigma+a\sigma^{\prime}$ or $\sigma_{\pm}$ Linear $\sigma$ centrally, ${\sigma^{+}}$ and ${\sigma^{-}}$ above and below. Discontinuous $L^{\prime}$
Edgeworth expansion Log of Equation (A.8) Numerically messy
Skew normal Log of Equation (42) Numerically messy
@@
As an illustration we take the result $a={5.0000}^{+2.5811}_{-1.9159}$, which would be reported from a Poisson measurement of 5 events. This is shown in Figure 10. The true likelihood is $L(a)=e^{-a}{a^{5}\over 5!}$, and we compare this with the results of the likelihoods given by the models. All curves, by construction, peak at $a=5$ and go through the $\Delta\ln L=\pm{1\over 2}$ points, and their interpolations in the central region are very similar. At larger values their behaviours are very different. The cubic turns over, which is unacceptable. The quartic provides a better match to the original Poisson likelihood, especially on the negative side. The logarithmic form does fairly well, the Poisson does perfectly, as it should in this particular example. The two Bartlett forms, labelled Linear Variance and Linear Sigma,both do well, the linear variance is somewhat better than the linear sigma. The PDG form does badly, as it assumes a Gaussian behaviour outside the central interval. The Edgeworth form does badly at both small and large values, and is troubled by logarithms of negative numbers at large values and/or asymmetries. The skew normal is not as bad as Edgeworth’s approximation, but it does not do very well. These comparisons assume the fundamental $a={5.0000}^{+2.5811}_{-1.9159}$ measurement comes from a Poisson measurement: if it does not, then conclusions might be different.
@@
As another example, this time with a negative skewness, we take the logarithm of a parameter with a likelihood described by a Gaussian with $\mu=8.0$ and $\sigma=3.0$. This is shown in Figure 11. The reported result is $a={2.08}^{+0.32}_{-0.47}$. Again the cubic does predictably badly, and the quartic surprisingly well. The logarithmic does quite well, and the Poisson is not good. The PDG form is again bad, but the Bartlett forms are good: this time the linear sigma does a bit better than the linear variance. The Edgeworth form does badly and the skew normal form does not do particularly well.
Figure 12: The central regions for Figures 10 and 11.